OpenStudy (anonymous):

Suppose that y varies inversely with x, and y = 6 when x = 8. What is an equation for the inverse variation?

OpenStudy (aroub):

yx=k

OpenStudy (anonymous):

y=k/x. Plug your values. 6=k/8, so k=48. Answer: y=48/x

OpenStudy (anonymous):

thank you